trentV1:

Your boss hands you a memo with a summary of the monthly data. The number of imports is shown as f(x), and the number of exports is shown as g(x). Use the data in the table below, representing both functions, to explain to your boss the solution to the system of equations and what that solution represents. Use complete sentences. Month f(x) = No. of imports g(x) = No. of exports January (1) 2 3 February (2) 4 4 March (3) 6 5 April (4) 8 6

Hero:

Provide a screenshot of the question from the book

trentV1:

1 attachment
Hero:

Okay, so you just have to figure out the equations for \(f(x)\) and \(g(x)\)

trentV1:

ok

Hero:

Any ideas on how to figure them out?

trentV1:

Not really im not to good at algebra

Hero:

Okay so the numbers under month are the x values correct?

trentV1:

Ok

Hero:

We have to figure out the relationship between \(x\) and \(f(x)\)

Hero:

And the relationship between \(x\) and \(g(x)\)

Hero:

So first \(x\) and \(f(x)\)

Hero:

\(\displaystyle \begin{array}{|c|c|} \hline x & f( x)\\ \hline 1 & 2\\ \hline 2 & 4\\ \hline 3 & 6\\ \hline 4 & 8\\ \hline \end{array}\)

Hero:

Notice that when \(x\) is \(1\), \(f(x)\) is \(2\) When \(x\) is \(2\) , \(f(x)\) is 4, and so on. Do you see a pattern there?

trentV1:

yea its going up by 2

Hero:

Actually, the correct thought to use in this case is \(f(x)\) DOUBLES the value of \(x\)

Hero:

Or rather \(f(x)\) is DOUBLE the value of \(x\)

Hero:

How could we write the expression for \(f(x)\) algebraically knowing this?

trentV1:

This part im kinda bad at but f(x)=1*2 idk

Hero:

Basically, from the table you observed that \(f(x)\) is TWICE the value of \(x\). How can we write the expression for \(x\) knowing this? Hint: The expression must be written in terms of \(x\).

trentV1:

ig f(2)=x

Hero:

\(f(x) = 2x\) actually

Hero:

Try finding the equation for \(g(x)\) now.

trentV1:

g(x)=1x

Hero:

What is the relationship between \(x\) and \(g(x)\) ?

trentV1:

+1

Hero:

Yes \(g(x)\) increases \(x\) by one.

Hero:

So what is the expression for \(g(x)\). Try writing it in terms of \(x\)

trentV1:

g(x)=1x

Hero:

How do you show that \(x\) is INCREASED by one?

Hero:

Your function shows that \(x\) is MULTIPLIED by \(1\) but not INCREASED by one

trentV1:

g(x)=x^1 or g(x)=x+1

trentV1:

I think i got this now i had my parents also help but i appreciate the help thank you for teaching me some of this.

Hero:

You're welcome